If 0.75 g of a monoprotic weak acid required 22.50 mL of 0.510 M NaOH to titrate it, what is the molar mass of the acid?

Select one:

33 g/mol

153 g/mol

51 g/mol

65 g/mol

86 g/mol
Answer #1

Balanced chemical equation is:
HA + NaOH ---> NaA + H2O

lets calculate the mol of HA
volume , V = 22.5 mL
= 2.25*10^-2 L


use:
number of mol,
n = Molarity * Volume
= 0.51*2.25*10^-2
= 1.147*10^-2 mol
According to balanced equation
mol of NaOH reacted = (1/1)* moles of HA
= (1/1)*1.147*10^-2
= 1.147*10^-2 mol
This is number of moles of NaOH

mass(NaOH)= 0.75 g


use:
number of mol = mass / molar mass
1.147*10^-2 mol = (0.75 g)/molar mass
molar mass = 65.36 g/mol
Answer: 65 g/mol